Jan needs 7 milliliters of a 55% solution. She only has 40% of weak solution and 61% of a strong solution. How many milliliters of each does she need of each solution.
Answer by mananth(16946)

Percent ---------------- quantity
Solution I 40.00% ---------------- x ml
Solution II 61.00% ------ 7 - x ml
Mixture 55.00% ---------------- 7
Total 7 ml
40.00% x + 61.00% ( 7 - x ) = 55.00% * 7
40 x + 61 ( 7 - x ) = 385
40 x + 427 - 61 x = 385
40 x - 61 x = 385 - -427
-21 x = -42
/ -21
x = 2 ml 40.00% Solution I
5 ml 61.00% Solution II

m.ananth@hotmail.ca